The vapor pressure of water at 80°C is 0.4675 atm. Calculate the vapor pressure in kPa. Be sure your answer has the right number of significant digits.


1
Expert's answer
2021-05-06T07:39:29-0400

"0.4675\ atm= 0.4675\times (1\ atm)"


"0.4675\ atm= 0.4675\times (1.01325\ bar)"


"0.4675\ atm\approx \boldsymbol{0.4737\ bar}"


And,


"0.4675\ atm=0.4737\times (1\ bar)"


"0.4675\ atm=0.4737\times (10^5\ Pa)"


"0.4675\ atm=\boldsymbol{47.37\ kPa}"


=47.37 kPa